OpenStudy (anonymous):

2c+6 over4-c * 4c-16 over 5c +15

OpenStudy (anonymous):

I have more questions after this if you stay on i"ll give you a medal

OpenStudy (anonymous):

\[\frac{ 2c+6 }{ 4c }\times \frac{ 4c-16 }{ 5c+16 }\]

OpenStudy (anonymous):

Like that?

OpenStudy (anonymous):

yes

OpenStudy (cgreenwade2000):

Factor those expressions!

OpenStudy (anonymous):

i dont know how!!!

OpenStudy (cgreenwade2000):

Okay, no problem. 2c+ 6 can be factored. Both terms are actually being multiplied by the same number (2). So we can factor out a 2 and get the following (2)(c+3) I didn't change anything about the expression besides what it looks like.

OpenStudy (cgreenwade2000):

Can you factor anything out out 4c-16 and 5c+15

OpenStudy (anonymous):

one sec